# Co-rewarding: Stable Self-supervised RL for Eliciting Reasoning in Large Language Models

### v2

@@ Abstract
While reinforcement learning with verifiable rewards (RLVR) is effective to improve the reasoning ability of large language models (LLMs), its reliance on human-annotated labels leads to the scaling up dilemma, especially for complex tasks. Recent self-rewarding methods investigate a label-free alternative to unlock the reasoning capabilities of LLMs, yet they frequently encounter the non-negligible training collapse issue, as the single-view supervision signal easily forms the self-consistent illusion, yielding the reward hacking. Inspired by the success of self-supervised learning, we propose Co-rewarding, a novel self-supervised RL framework that improves training stability by seeking complementary supervision from another views. Specifically, we instantiate Co-rewarding in two ways: (1) Co-rewarding-I is a data-side instantiation that derives reward signals from contrastive agreement across semantically analogous questions; and (2) Co-rewarding-II is a model-side instantiation that maintains a slowly-updated reference teacher with pseudo labels to realize self-distillation. Intuitively, such instantiations introduce different levels of discrepancy to increase the difficulty of training collapse on trivial reasoning solutions. Empirically, Co-rewarding exhibits stable training across various setups, and outperforms other self-rewarding baselines by $3.31\%$ improvements on average on multiple mathematical reasoning benchmarks, especially by $7.49\%$ on Llama-3.2-3B-Instruct. Notably, Co-rewarding reaches or even surpasses RLVR with ground-truth (GT) label in several cases, such as a Pass@$1$ of $94.01\%$ on GSM8K with Qwen3-8B-Base remarkably higher than GT. Our code is publicly available at https://github.com/tmlr-group/Co-rewarding.
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024; Qwen et al., 2025) have demonstrated remarkable general-purpose capabilities in a wide range of linguistic tasks (Hendrycks et al., ). To further elicit their reasoning ability in complex scenarios, reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; Yu et al., 2025) is developed for post-training with externally verifiable signals like program execution results (Luo et al., 2025) or mathematical equivalence (Shao et al., 2024). Despite the impressive improvement, the reliance on high-quality ground-truth (GT) labels of RLVR remains as a major bottleneck (Ouyang et al., 2022; Bai et al., 2022) in the spirit of the scaling law, which subsequently motivates the emerging exploration of self-rewarding methods with unlabeled data (Zhao et al., 2025b; Zuo et al., 2025; Zhang et al., 2025b).
@@
One prominent line of such label-free methods leverages the internal signals (e.g., entropy (Zhang et al., 2025c; Prabhudesai et al., 2025) and self-certainty (Zhao et al., 2025b)) to strengthen the confidence of the model in reasoning. Another critical line seeks the answer-level consensus (Zuo et al., 2025; Shafayat et al., 2025) to construct pseudo labels as reward basis. while effective initially, these self-rewarding approaches frequently exhibit non-negligible training collapse (Zhang et al., 2025d) (indicated as right of Figure 1), which limits the scalability of such label-free training manners.
@@
The collapse phenomenon stems from reward hacking (Laidlaw et al., 2025) under self-consistent illusion: the reward signal is internally produced by the policy model from a single-view data perspective, which is easily trapped by trivial solutions along with training (see Figure 7). Specifically, for entropy- or certainty-based rewards, the policy model may concentrate probability mass on a small set of tokens and produce repetitive strings that minimize entropy or maximize self-certainty (Zhang et al., 2025d). And for consensus-based rewards, the policy model can converge to a consistent yet incorrect answer that attains high consensus across rollouts (Shafayat et al., 2025). Overall, the policy model continually reduces uncertainty without sustained gains in correctness, inflating the reward but eroding exploration and diversity. It ultimately collapses when a persistent hacking strategy emerges.
@@
To this end, we introduce Co-rewarding, a self-supervised RL framework that seeks complementary supervision from another views, inspired by self-supervised learning (Chen et al., 2020; Grill et al., 2020; Caron et al., 2021). Conceptually, one fundamental characteristic of self-rewarding methods lies on that supervision intertwined with current policy on single-view outputs, for which we propose to seek reasoning invariance across different views (see Figure 2). Specifically, we investigate two initiations of Co-rewarding: (1) Co-rewarding-I: a data-side initiation that constructs rewards via contrastive agreement across semantically analogous questions, each providing pseudo labels for the other; and (2) Co-rewarding-II: a model-side initiation that introduces an extra teacher with dynamically updated policy and provides stable pseudo-labels insulated from current online policy.
@@ 3.1 Conceptual Philosophy: invariance beyond the single-view
This philosophy yields our Co-rewarding framework, whose core idea is to ground self-supervised RL in invariance rather than the suspicious single-view feedback. We instantiate it in two orthogonal ways: by enforcing analogy-invariance on the data side (Co-rewarding-I) and by disentangling supervision through temporal invariance on the model side (Co-rewarding-II).
@@ 3.2 Two Initiations of Co-rewarding Framework
Remark 1. Overall, the two instantiations of Co-rewarding embody our core idea from different perspectives: I leverages data-side analogy-invariance to provide cross supervision, while II employs model-side self-distillation to stabilize learning. Together, they reflect that stable self-supervised reasoning elicitation can emerge from both the diversity of data perspectives and the disentanglement of supervision signals. Moreover, Co-rewarding offers a flexible framework, in which key components, such as pseudo-labeling strategies, data rephrasing techniques, teacher model update rules, and policy optimization, can be seamlessly substituted with other advanced approaches (Yu et al., 2025).
@@ 4.1 Setups
Implementation Details. We implement our algorithms based on the VeRL framework (Sheng et al., 2024), and experiments are conducted on 4 $\times$ H100-80GB GPUs. For our experiments, we totally use three training sets: MATH (Hendrycks et al., 2021) (7,500 questions), DAPO-14k (Yu et al., 2025) (en-version of DAPO-Math-17k, about 14.1k questions), and OpenRS (Dang & Ngo, 2025) (7,000 questions). During RL training, we use a global batch size of 128, set the number of rollouts to $G=\tilde{G}=8$ per question for both I and II, and adopt AdamW with a learning rate of $3\times 10^{-6}$. In Co-rewarding-I, question rephrasing is performed by the open-source Qwen3-32B model. In Co-rewarding-II, the EMA weight is scheduled from $\alpha_{\textrm{start}}=0.99$ to $\alpha_{\textrm{end}}=0.9999$ using cosine annealing. More implementation details are reported in Appendix C.2.
Training Set: MATH Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
@@
Before RL 72.4 27.82 20.93 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 82.6 87.26 54.22 30.52 63.25 52.78 57.11
- Self-Certainty (Zhao et al., 2025b) 80.2 80.74 50.75 27.20 64.38 50.98 54.17
- Entropy (Prabhudesai et al., 2025) 80.2 87.19 49.54 29.38 62.00 51.81 54.86
- Majority-Voting (Shafayat et al., 2025) 79.8 89.76 49.09 30.52 63.38 51.80 56.93
- Co-rewarding-I (Ours) 81.2 93.70 51.20 30.81 66.00 55.79 59.95
- Co-rewarding-II (Ours) 80.8 92.42 53.46 30.23 62.83 60.70 57.50
@@
Before RL 71.2 26.15 21.08 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 78.6 89.76 51.20 26.07 55.38 47.80 53.96
- Self-Certainty (Zhao et al., 2025b) 71.6 71.79 38.86 22.37 57.00 48.15 48.93
- Entropy (Prabhudesai et al., 2025) 77.0 88.10 47.44 25.59 52.88 50.44 49.90
- Majority-Voting (Shafayat et al., 2025) 77.4 90.07 45.33 26.54 57.50 48.78 54.35
- Co-rewarding-I (Ours) 78.8 91.28 46.08 26.64 56.50 50.35 53.26
- Co-rewarding-II (Ours) 78.0 88.86 45.93 26.25 55.00 51.30 53.88
@@
Before RL 39.2 65.73 10.54 9.86 25.37 57.32 31.14
- GT-Reward (Shao et al., 2024) 47.0 77.94 22.14 9.57 31.87 47.51 34.32
- Self-Certainty (Zhao et al., 2025b) 43.4 74.91 18.83 9.95 25.87 54.88 33.34
- Entropy (Prabhudesai et al., 2025) 43.4 66.19 20.18 11.66 24.62 54.70 33.52
- Majority-Voting (Shafayat et al., 2025) 46.8 78.77 20.48 11.00 31.25 47.96 33.18
- Co-rewarding-I (Ours) 50.2 79.45 23.80 11.28 29.88 48.89 33.77
- Co-rewarding-II (Ours) 49.8 79.30 22.59 10.80 30.63 49.90 33.61
Evaluation Details. To provide a comprehensive evaluation of model capabilities, we utilize a diverse set of benchmarks spanning mathematical reasoning, code generation, instruction-following, and general multi-task abilities. Specifically: (1) Mathematical reasoning: MATH500 (Lightman et al., 2024), GSM8K (Cobbe et al., 2021), and AMC (Li et al., 2024a). (2) Code generation: LiveCodeBench (Jain et al., 2025) release_v6 and CRUX (Gu et al., 2024). (3) Instruction-following and multi-task abilities: IFEval (Zhou et al., 2023b) and MMLU-Pro (Wang et al., 2024). Additional evaluation details are provided in Appendix C.3.
@@ 4.2.1 Main Performance of Co-rewarding
Superior Performance of Co-rewarding over self-rewarding baselines. Table 1 and Table 2 report the experimental results trained on MATH and DAPO-14k, respectively. We observe that both Co-rewarding-I and Co-rewarding-II occupy more darker cells in the tables, demonstrating stronger performance than other self-rewarding SoTA baselines. Specifically, Co-rewarding-I achieves an average relative performance gain of $3.46\%$ over the best baselines across three mathematical benchmarks and models in Table 1, while Co-rewarding-II achieves a larger average relative gain of $7.29\%$ in Table 2. Additional results on other training sets and LLMs are provided in Appendix D.1.
Surpassing GT-Reward in certain benchmarks. Surprisingly, we observe that both Co-rewarding-I and Co-rewarding-II outperform GT-Reward in certain cases. On GSM8K, they together achieve an average relative improvement of $2.94\%$ over GT-Reward in Table 1, while Co-rewarding-II further delivers a larger gain of $5.44\%$ in Table 2. Co-rewarding also shows advantages on the coding benchmark CRUX in several cases. Notably, Co-rewarding-II achieves a remarkably high Pass@1 of 94.01% with Qwen3-8B-Base on GSM8K. These findings highlight the potential of self-supervised RL to elicit stronger reasoning abilities, particularly with Co-rewarding mitigating training collapse.
Table 2: Main Results (%) of Co-rewarding and baselines trained on DAPO-14k. Cell background colors indicate relative performance: darker colors denote better results within each model group. Additional Results of Qwen3-8B-Base trained on OpenRS refer to Table 8.
Training Set: DAPO-14k Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Before RL 72.4 27.82 20.93 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 86.6 87.19 61.75 30.52 63.75 53.11 60.27
- Self-Certainty (Zhao et al., 2025b) 82.0 77.63 49.85 27.77 60.75 50.58 54.24
- Entropy (Prabhudesai et al., 2025) 79.4 80.82 45.48 30.14 62.00 51.56 54.57
- Majority-Voting (Shafayat et al., 2025) 78.6 91.66 50.00 30.33 61.62 51.54 55.65
- Co-rewarding-I (Ours) 78.4 88.02 51.20 29.38 62.50 50.17 55.39
- Co-rewarding-II (Ours) 80.6 94.01 54.37 31.66 67.12 53.31 59.83
Qwen3-4B-Base
Before RL 71.2 26.15 21.08 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 83.6 85.14 52.86 18.58 56.88 47.70 55.35
- Self-Certainty (Zhao et al., 2025b) 68.4 44.81 35.39 25.88 50.12 45.58 48.84
- Entropy (Prabhudesai et al., 2025) 76.6 82.79 43.37 26.35 50.75 48.20 50.22
- Majority-Voting (Shafayat et al., 2025) 73.4 64.06 40.81 26.16 53.00 48.91 51.06
- Co-rewarding-I (Ours) 73.8 75.89 43.83 26.25 50.12 46.84 51.51
- Co-rewarding-II (Ours) 77.8 91.89 48.49 26.64 54.87 48.90 52.83
Llama-3.2-3B-Instruct
Before RL 39.2 65.73 10.54 9.86 25.37 57.32 31.14
- GT-Reward (Shao et al., 2024) 49.4 78.17 25.90 10.33 31.37 53.10 33.83
- Self-Certainty (Zhao et al., 2025b) 42.4 74.71 17.32 11.18 28.38 54.50 33.51
- Entropy (Prabhudesai et al., 2025) 44.0 65.85 17.32 9.95 25.00 55.78 31.95
- Majority-Voting (Shafayat et al., 2025) 42.8 70.96 17.62 10.14 29.50 54.07 32.95
- Co-rewarding-I (Ours) 46.0 70.58 20.93 9.57 27.25 53.04 32.61
- Co-rewarding-II (Ours) 49.8 78.62 19.73 10.43 32.25 51.92 34.46
Code generalization with preserved general performance. Although trained solely on math-oriented datasets, the models show improvements on coding benchmarks, suggesting a cross-domain generalization from math to code in self-supervised reasoning elicitation. Moreover, Co-rewarding preserves general instruction-following and multi-task ability on MMLU-Pro and IFEval. As shown in Table 4, Co-rewarding-II outperforms other self-rewarding baselines in 12 of 14 MMLU-Pro categories, demonstrating that its gains do not come at the expense of broader general-domain performance. More detailed results of MMLU-Pro and IFEval refer to Appendix D.7 and D.8.
@@
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Co-rewarding-I 81.2 93.70 51.20 30.81 66.00 55.79 59.95
- Majority-Voting w/ Original 79.8 89.76 49.09 30.52 63.38 51.80 56.93
- Majority-Voting w/ Rephrased 79.2 91.51 50.75 31.66 60.38 52.24 57.26
Co-rewarding-II 80.6 94.01 54.37 31.66 67.12 53.31 59.83
- w/o Updating Reference 78.0 88.40 51.66 30.62 63.75 52.48 58.01
@@
Co-rewarding-I 50.2 79.45 23.80 11.28 29.88 48.89 33.77
- Majority-Voting w/ Original 46.8 78.77 20.48 11.00 31.25 47.96 33.18
- Majority-Voting w/ Rephrased 44.0 78.85 21.23 10.04 17.25 47.84 33.72
Co-rewarding-II 49.8 78.62 19.73 10.43 32.25 51.92 34.46
- w/o Updating Reference 45.0 76.72 17.92 10.05 30.63 51.33 33.94
4.2.2 Further Analysis
Co-rewarding alleviates collapse and provides stable self-supervised RL. We use 5,000 questions from the MATH test split as a validation set to monitor training process. Figure 4 shows the all three self-rewarding baselines collapse on both MATH and DAPO-14k. Co-rewarding-I remains stable on MATH but still collapses on DAPO-14k, suggesting that its stability depends on the property of training data. A plausible explanation is that MATH contains more questions with rich background descriptions than DAPO-14k, which makes rephrasing more effective in amplifying differences between original and rephrased data, thereby strengthening contrastive agreement. In contrast, Co-rewarding-II consistently maintains stability across datasets, as its design decouples supervision from the online policy and thus breaks the entanglement between supervision signals and the policy itself.
@@ 4.2.2 Further Analysis
Figure 6: Pseudo label accuracy comparison of Co-rewarding-II, “w/o updating reference”, and Majority-Voting.

### v3

@@ Abstract
While reinforcement learning with verifiable rewards (RLVR) is effective to improve the reasoning ability of large language models (LLMs), its reliance on human-annotated labels leads to the scaling up dilemma, especially for complex tasks. Recent self-rewarding methods investigate a label-free alternative to unlock the reasoning capabilities of LLMs, yet they frequently encounter the non-negligible training collapse issue, as the single-view supervision signal easily forms the self-consistent illusion, yielding the reward hacking. Inspired by the success of self-supervised learning, we propose Co-rewarding, a novel self-supervised RL framework that improves training stability by seeking complementary supervision from another views. Specifically, we instantiate Co-rewarding in two ways: (1) Co-rewarding-I is a data-side instantiation that derives reward signals from contrastive agreement across semantically analogous questions; and (2) Co-rewarding-II is a model-side instantiation that maintains a slowly-updated reference teacher with pseudo labels to realize self-distillation. Intuitively, such instantiations introduce different levels of discrepancy to increase the difficulty of training collapse on trivial reasoning solutions. We also explore their orthogonally combined version to further boost the performance. Empirically, Co-rewarding exhibits stable training across various setups, and outperforms other self-rewarding baselines by $3.31\%$ improvements on average on multiple mathematical reasoning benchmarks, especially by $7.49\%$ on Llama-3.2-3B-Instruct. Notably, Co-rewarding reaches or even surpasses RLVR with ground-truth (GT) label in several cases, such as a Pass@$1$ of $94.01\%$ on GSM8K with Qwen3-8B-Base remarkably higher than GT. Our code is released at https://github.com/tmlr-group/Co-rewarding.
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024) have demonstrated remarkable general-purpose capabilities in a wide range of linguistic tasks (Li et al., 2026; Chen et al., 2026). To further elicit their reasoning ability in complex scenarios, reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; Yu et al., 2025) is developed for post-training with externally verifiable signals like program execution results (Luo et al., 2025) or mathematical equivalence (Shao et al., 2024). Despite the impressive improvement, the reliance on high-quality ground-truth (GT) labels of RLVR remains as a major bottleneck (Ouyang et al., 2022; Bai et al., 2022) in the spirit of the scaling law, which subsequently motivates the emerging exploration of self-rewarding methods with unlabeled data (Zhao et al., 2025b; Zuo et al., 2025; Zhang et al., 2025c).
@@
One prominent line of such label-free methods leverages the internal signals (e.g., entropy (Zhang et al., 2025d; Prabhudesai et al., 2025) and self-certainty (Zhao et al., 2025b)) to strengthen the confidence of the model in reasoning. Another critical line seeks the answer-level consensus (Zuo et al., 2025; Shafayat et al., 2025) to construct pseudo labels as reward basis. while effective initially, these self-rewarding approaches frequently exhibit non-negligible training collapse (Zhang et al., 2025e) (indicated as right of Figure 1), which limits the scalability of such label-free training manners.
@@
The collapse phenomenon stems from reward hacking (Laidlaw et al., 2025) under self-consistent illusion: the reward signal is internally produced by the policy model from a single-view data perspective, which is easily trapped by trivial solutions along with training (see Figure 7). Specifically, for entropy- or certainty-based rewards, the policy model may concentrate probability mass on a small set of tokens and produce repetitive strings that minimize entropy or maximize self-certainty (Zhang et al., 2025e). And for consensus-based rewards, the policy model can converge to a consistent yet incorrect answer that attains high consensus across rollouts (Shafayat et al., 2025). Overall, the policy model continually reduces uncertainty without sustained gains in correctness, inflating the reward but eroding exploration and diversity. It ultimately collapses when a persistent hacking strategy emerges.
@@
To this end, we introduce Co-rewarding, a self-supervised RL framework that seeks complementary supervision from another views, inspired by self-supervised learning (Chen et al., 2020; Grill et al., 2020; Caron et al., 2021). Conceptually, one fundamental characteristic of self-rewarding methods lies on that supervision intertwined with current policy on single-view outputs, for which we propose to seek reasoning invariance across different views (see Figure 2). Specifically, we investigate two initiations of Co-rewarding: (1) Co-rewarding-I: a data-side initiation that constructs rewards via contrastive agreement across semantically analogous questions, each providing pseudo labels for the other; and (2) Co-rewarding-II: a model-side initiation that introduces an extra teacher with dynamically updated policy and provides stable pseudo-labels insulated from current online policy. Additionally, we also explore the combined instantiation, Co-rewarding-III, which integrates data-side cross-supervision with model-side self-distillation to further boost the performance.
@@ 3.1 Conceptual Philosophy: invariance beyond the single-view
This philosophy yields our Co-rewarding framework, whose core idea is to ground self-supervised RL in invariance rather than the suspicious single-view feedback.We instantiate it in two orthogonal ways and one combined version: by enforcing analogy-invariance on the data side (Co-rewarding-I), by disentangling supervision through temporal invariance on the model side (Co-rewarding-II), and by integrating both mechanisms in a unified instantiation (Co-rewarding-III).
@@ 3.2 Two Initiations of Co-rewarding Framework
Co-rewarding-III: Data-side + Model-side. Given that Co-rewarding-I and Co-rewarding-II provide two complementary perspectives for constructing stable self-supervised signals, a natural exploration is to integrate both data-side cross-supervision and model-side self-distillation into a unified instantiation. We introduce Co-rewarding-III, which leverages analogy-invariance between each original question and its rephrased counterparts while producing pseudo-labels from the EMA-updated reference teacher. Specifically, the teacher generates rollouts for both original and rephrased questions, and the resulting pseudo-label from one side is used to supervise the other. This combination further boosts the resistance of the reward signal to hacking, promoting more stable training dynamics.
Formally, its learning objective can be formulated as:
[…]
where the first term supervises the original question via pseudo labels generated from its rephrased counterpart, and the second term, symmetrically, supervises the rephrased question via pseudo labels generated from the original question. The estimated advantages $\mathcal{R}_{\theta}(\hat{A}^{(k)})$ and $\mathcal{R}_{\theta}(\hat{A}^{\prime(k)})$ are computed in the similar way as in Co-rewarding-I and Co-rewarding-II. The reference teacher is also updated via EMA, as Eq. (11) in Co-rewarding-II. The other formulations and pseudo code of Co-rewarding-III are supplemented in Appendix B.1 and Algorithm 3.
Remark 1. Overall, the two instantiations of Co-rewarding embody our core idea from different perspectives: I leverages data-side analogy-invariance to provide cross supervision, while II employs model-side self-distillation to stabilize learning. Together, they reflect that stable self-supervised reasoning elicitation can emerge from both the diversity of data perspectives and the disentanglement of supervision signals. Co-rewarding-III further explores an orthogonally combined instantiation of these two sides. Moreover, Co-rewarding offers a flexible framework, in which key components, such as pseudo-labeling strategies, data rephrasing techniques, teacher model update rules, and policy optimization, can be seamlessly substituted with other advanced approaches (Yu et al., 2025).
Training Set: MATH Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC AIME24 LiveCode CRUX IFEval MMLU-Pro
Qwen3-8B-Base
Before RL 72.4 27.82 20.93 3.75 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 82.6 87.26 54.22 17.15 30.52 63.25 52.78 57.11
- Self-Certainty (Zhao et al., 2025b) 80.2 80.74 50.75 15.73 27.20 64.38 50.98 54.17
- Entropy (Prabhudesai et al., 2025) 80.2 87.19 49.54 15.63 29.38 62.00 51.81 54.86
- Majority-Voting (Shafayat et al., 2025) 79.8 89.76 49.09 15.83 30.52 63.38 51.80 56.93
- Co-rewarding-I (Ours) 81.2 93.70 51.20 15.10 30.81 66.00 55.79 59.95
- Co-rewarding-II (Ours) 80.8 92.42 53.46 14.48 30.23 62.83 60.70 57.50
- Co-rewarding-III (Ours) 81.4 90.98 54.07 13.33 30.71 63.75 53.69 59.10
Qwen3-4B-Base
Before RL 71.2 26.15 21.08 4.58 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 78.6 89.76 51.20 15.00 26.07 55.38 47.80 53.96
- Self-Certainty (Zhao et al., 2025b) 71.6 71.79 38.86 11.67 22.37 57.00 48.15 48.93
- Entropy (Prabhudesai et al., 2025) 77.0 88.10 47.44 10.94 25.59 52.88 50.44 49.90
- Majority-Voting (Shafayat et al., 2025) 77.4 90.07 45.33 10.10 26.54 57.50 48.78 54.35
- Co-rewarding-I (Ours) 78.8 91.28 46.08 13.85 26.64 56.50 50.35 53.26
- Co-rewarding-II (Ours) 78.0 88.86 45.93 12.17 26.25 55.00 51.30 53.88
- Co-rewarding-III (Ours) 78.6 90.75 48.80 12.71 26.16 56.00 49.23 53.08
Llama-3.2-3B-Instruct
Before RL 39.2 65.73 10.54 3.75 9.86 25.37 57.32 31.14
- GT-Reward (Shao et al., 2024) 47.0 77.94 22.14 11.67 9.57 31.87 47.51 34.32
- Self-Certainty (Zhao et al., 2025b) 43.4 74.91 18.83 6.88 9.95 25.87 54.88 33.34
- Entropy (Prabhudesai et al., 2025) 43.4 66.19 20.18 6.56 11.66 24.62 54.70 33.52
- Majority-Voting (Shafayat et al., 2025) 46.8 78.77 20.48 9.27 11.00 31.25 47.96 33.18
- Co-rewarding-I (Ours) 50.2 79.45 23.80 10.00 11.28 29.88 48.89 33.77
- Co-rewarding-II (Ours) 49.8 79.30 22.59 10.73 10.80 30.63 49.90 33.61
- Co-rewarding-III (Ours) 51.6 79.91 25.45 10.42 10.43 32.50 46.37 34.50
@@ 4.1 Setups
Table 2: Main Results (%) of Co-rewarding and baselines trained on DAPO-14k. Cell background colors indicate relative performance: darker colors denote better results within each model group. Additional Results of Qwen3-8B-Base and Qwen3-4B-Base trained on OpenRS refer to Table 8.
Training Set: DAPO-14k Mathematics Code Instruction Multi-Task
Methods MATH500 GSM8K AMC AIME24 LiveCode CRUX IFEval MMLU-Pro
@@
Before RL 72.4 27.82 20.93 3.75 23.41 54.75 50.89 52.92
- GT-Reward (Shao et al., 2024) 86.6 87.19 61.75 24.58 30.52 63.75 53.11 60.27
- Self-Certainty (Zhao et al., 2025b) 82.0 77.63 49.85 15.00 27.77 60.75 50.58 54.24
- Entropy (Prabhudesai et al., 2025) 79.4 80.82 45.48 15.00 30.14 62.00 51.56 54.57
- Majority-Voting (Shafayat et al., 2025) 78.6 91.66 50.00 11.25 30.33 61.62 51.54 55.65
- Co-rewarding-I (Ours) 78.4 88.02 51.20 11.88 29.38 62.50 50.17 55.39
- Co-rewarding-II (Ours) 80.6 94.01 54.37 16.35 31.66 67.12 53.31 59.83
- Co-rewarding-III (Ours) 81.6 92.27 53.77 17.71 32.70 66.75 55.85 60.02
@@
Before RL 71.2 26.15 21.08 4.58 11.00 38.88 46.43 47.23
- GT-Reward (Shao et al., 2024) 83.6 85.14 52.86 20.63 18.58 56.88 47.70 55.35
- Self-Certainty (Zhao et al., 2025b) 68.4 44.81 35.39 8.85 25.88 50.12 45.58 48.84
- Entropy (Prabhudesai et al., 2025) 76.6 82.79 43.37 12.81 26.35 50.75 48.20 50.22
- Majority-Voting (Shafayat et al., 2025) 73.4 64.06 40.81 9.17 26.16 53.00 48.91 51.06
- Co-rewarding-I (Ours) 73.8 75.89 43.83 10.63 26.25 50.12 46.84 51.51
- Co-rewarding-II (Ours) 77.8 91.89 48.49 14.27 26.64 54.87 48.90 52.83
- Co-rewarding-III (Ours) 79.2 90.45 48.95 15.10 27.58 54.87 50.30 54.79
@@
Before RL 39.2 65.73 10.54 3.75 9.86 25.37 57.32 31.14
- GT-Reward (Shao et al., 2024) 49.4 78.17 25.90 9.17 10.33 31.37 53.10 33.83
- Self-Certainty (Zhao et al., 2025b) 42.4 74.71 17.32 4.79 11.18 28.38 54.50 33.51
- Entropy (Prabhudesai et al., 2025) 44.0 65.85 17.32 6.56 9.95 25.00 55.78 31.95
- Majority-Voting (Shafayat et al., 2025) 42.8 70.96 17.62 8.74 10.14 29.50 54.07 32.95
- Co-rewarding-I (Ours) 46.0 70.58 20.93 7.08 9.57 27.25 53.04 32.61
- Co-rewarding-II (Ours) 49.8 78.62 19.73 8.02 10.43 32.25 51.92 34.46
- Co-rewarding-III (Ours) 48.6 76.95 21.84 8.13 9.86 30.50 49.92 34.01
Implementation Details. We implement our algorithms based on the VeRL framework (Sheng et al., 2024), and experiments are conducted on 4 $\times$ H100-80GB GPUs. For our experiments, we totally use three training sets: MATH (Hendrycks et al., 2021) (7,500 questions), DAPO-14k (Yu et al., 2025) (en-version of DAPO-Math-17k, about 14.1k questions), and OpenRS (Dang and Ngo, 2025) (7,000 questions). During RL training, we use a global batch size of 128, set the number of rollouts to $G=\tilde{G}=8$ per question for Co-rewarding-I, II and III, and adopt AdamW with a learning rate of $3\times 10^{-6}$. In Co-rewarding-I and III, question rephrasing is performed by the open-source Qwen3-32B model. In Co-rewarding-II and III, the EMA weight is scheduled from $\alpha_{\textrm{start}}=0.99$ to $\alpha_{\textrm{end}}=0.9999$ using cosine annealing. More implementation details are reported in Appendix C.2.
Evaluation Details. To provide a comprehensive evaluation of model capabilities, we utilize a diverse set of benchmarks spanning mathematical reasoning, code generation, instruction-following, and general multi-task abilities. Specifically: (1) Mathematical reasoning: MATH500 (Lightman et al., 2024), GSM8K (Cobbe et al., 2021), AMC (Li et al., 2024a), and AIME24 (Zhang and Math-AI, 2024). (2) Code generation: LiveCodeBench (Jain et al., 2025) release_v6 and CRUX (Gu et al., 2024). (3) Instruction-following and multi-task abilities: IFEval (Zhou et al., 2023b) and MMLU-Pro (Wang et al., 2024). Additional evaluation details are provided in Appendix C.3.
@@ 4.2.1 Main Performance of Co-rewarding
Superior Performance of Co-rewarding over self-rewarding baselines. Table 1 and Table 2 report the experimental results trained on MATH and DAPO-14k, respectively. We observe that all three Co-rewarding instantiations (I, II, and III) occupy more darker cells in the tables, demonstrating stronger performance than other self-rewarding SoTA baselines. Specifically, Co-rewarding-I achieves an average relative performance gain of $4.42\%$ over the best baselines across four mathematical benchmarks and models in Table 1, while Co-rewarding-II achieves a larger average relative gain of $12.90\%$ in Table 2. Moreover, Co-rewarding-III achieves improvements on average of $7.11\%$ and $1.72\%$ over Co-rewarding-I and Co-rewarding-II, respectively, suggesting that integrating data-side cross-supervision with model-side self-distillation can further boost performance. Additional results on other training sets and LLMs are provided in Appendix D.1.
Surpassing GT-Reward on certain benchmarks. Surprisingly, we observe that all three Co-rewarding instantiations (I, II, and III) outperform GT-Reward in certain cases. For example, on GSM8K, they together achieve an average improvement of $2.77\%$ over GT-Reward in Table 1, while Co-rewarding-II further delivers a larger gain of $5.44\%$ in Table 2. Notably, Co-rewarding-II reaches a remarkably high Pass@1 of 94.01% with Qwen3-8B-Base. This may be because GSM8K is a relatively easier benchmark, where self-supervised RL is sufficient to elicit latent reasoning abilities without GT labels. Additionally, Co-rewarding also shows advantages on coding benchmark CRUX in several cases, possibly due to distribution mismatch between training and evaluation data. This may offer opportunities for self-supervised methods to generalize on par with, or even surpass GT-supervised methods in some cases. These findings highlight the potential of self-supervised RL to elicit reasoning capabilities, particularly with Co-rewarding mitigating training collapse.
@@
Training Set Methods MATH500 GSM8K AMC AIME24 LiveCode CRUX IFEval MMLU-Pro
MATH Qwen3-8B-Base
Co-rewarding-I 81.2 93.70 51.20 15.10 30.81 66.00 55.79 59.95
- Majority-Voting w/ Union 80.2 93.48 49.70 15.63 31.94 64.88 54.25 59.51
- Majority-Voting w/ Original 79.8 89.76 49.09 15.83 30.52 63.38 51.80 56.93
- Majority-Voting w/ Rephrased 79.2 91.51 50.75 14.17 31.66 60.38 52.24 57.26
Co-rewarding-II 80.8 92.42 53.46 14.48 30.23 62.83 60.70 57.50
- w/o Updating Reference 79.2 89.46 51.51 13.96 30.62 61.75 56.93 51.85
@@
Co-rewarding-I 50.2 79.45 23.80 10.00 11.28 29.88 48.89 33.77
- Majority-Voting w/ Union 48.0 80.52 21.84 9.69 10.14 30.00 43.35 34.05
- Majority-Voting w/ Original 46.8 78.77 20.48 9.27 11.00 31.25 47.96 33.18
- Majority-Voting w/ Rephrased 44.0 78.85 21.23 8.85 10.04 17.25 47.84 33.72
Co-rewarding-II 49.8 79.30 22.59 10.73 10.80 30.63 49.90 33.61
- w/o Updating Reference 47.0 78.92 22.29 9.06 5.50 31.25 47.88 33.32
DAPO-14k Qwen3-8B-Base
Co-rewarding-II 80.6 94.01 54.37 16.35 31.66 67.12 53.31 59.83
- w/o Updating Reference 78.0 88.40 51.66 15.94 30.62 63.75 52.48 58.01
Llama-3.2-3B-Instruct
Co-rewarding-II 49.8 78.62 19.73 8.02 10.43 32.25 51.92 34.46
- w/o Updating Reference 45.0 76.72 17.92 8.02 10.05 30.63 51.33 33.94
Code generalization with preserved general performance. Although trained solely on math-oriented datasets, the models show improvements on coding benchmarks, suggesting a cross-domain generalization from math to code in self-supervised reasoning elicitation. Moreover, Co-rewarding preserves general instruction-following and multi-task ability on MMLU-Pro and IFEval. As shown in Table 4, Co-rewarding-II outperforms other self-rewarding baselines in 12 of 14 MMLU-Pro categories, demonstrating that its gains do not come at the expense of broader general-domain performance. More detailed results of MMLU-Pro and IFEval refer to Appendix D.10 and D.11.
@@
Figure 6: Pseudo label accuracy comparison.